OpenStudy (anonymous):

Differentiate: y=(3x+2)^cos(x)

OpenStudy (anonymous):

\[D\left[(2+3 x)^{\text{Cos}[x]},x\right]=(2+3 x)^{\text{Cos}[x]} \left(\frac{3 \text{Cos}[x]}{2+3 x}-\text{Log}[2+3 x] \text{Sin}[x]\right) \]